HWB4-DueApril26 (1)-Saved to my Mac Rohwedder and Willis (2010) gave memory tests to people aged 60 to 64 years in some European and North American countries. Their results suggested that early retirement may lead to memory decline. 7. The table shows some of their data (NOTE: a greater Percent Retired means that more people retired younger). Percent Retired Memory Score Country Sweden USA 39 48 59 70 74 78 81 87 9.3 10.9 10.7 9.1 6.4 9.1 7.2 England Germany Spain Netherlands Italy France Belgium Austria 7.9 8.5 9.0 าร 91 Use the four-step process to explain your analysis. Write up the results using APA format. a. b. MacBook Air
Answer #1

(a)

Hypotheses are:

H_{0}:\rho=0

H_{a}:\rho\neq 0

Following table shows the calculations:

X Y X^2 Y^2 XY
39 9.3 1521 86.49 362.7
48 10.9 2304 118.81 523.2
59 10.7 3481 114.49 631.3
70 9.1 4900 82.81 637
74 6.4 5476 40.96 473.6
78 9.1 6084 82.81 709.8
81 7.2 6561 51.84 583.2
87 7.9 7569 62.41 687.3
88 8.5 7744 72.25 748
91 9 8281 81 819
Total 715 88.1 53921 793.87 6175.1

Sample size: n=10

Now,

S_{yy}=\sum y^{2}-\frac{\left (\sum y \right )^{2}}{n}=17.709

S_{xx}=\sum x^{2}-\frac{\left (\sum x \right )^{2}}{n}=2798.5

S_{xy}=\sum xy-\frac{\left (\sum x \right )\left (\sum y \right )}{n}=-124.05

The coeffcient of correlation is :

r=\frac{S_{xy}}{\sqrt{S_{xx}S_{yy}}}=-0.557

Test statistics:

t=\frac{r}{\sqrt{\frac{1}{n-2}(1-r^{2})}}=-1.897

-------------------------

Here degree of freedom is df=n-2=8 so p-value of the test is 0.0944.

-----------------

Because the P-value is gereater than the significance level 0.05, there is not sufficient evidence to support the claim that there is a linear correlation between the variables.

b)

t(8) = -1.897, p > 0.05